Calculate the total number of atoms and total mass of 99mTc present in 15 mCi (555 MBq) 99mTc activity (t1/2 = 6 h).


1
Expert's answer
2021-10-13T09:08:31-0400

"A=\frac {N\ln2}{T_{1/2}},\implies"

"N=\frac{AT_{1/2}}{\ln 2}=17.3\cdot 10^{12};"

"N=\frac mMN_A,\implies"

"m=\frac{NM}{N_A}=2.85\cdot 10^{-12}~kg."